RC meeting and Run Plan, Monday February 13th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Report



Current Status

  • Running 45 uA on He3

Run Plan

  • Continue running 45 uA on He3
  • Accelerator maintenece Wednesday or Thursday

Data Collection and Polarization Status

  • 30.2C of uncorrected charge for this kinematic (34.9C of corrected charge)
  • Last NMR 48.7%

Other Topics/Status Updates

  • Anything else?
  • Bill wants to change the raster settings after some analysis

Requested Accesses to the Hall

  • Jack would like to collect some cables
  • GEM team need like a 3-4hour access for the SBS GEMs DAQ troubleshooting